# PASER [Uncaptioned image]: Post-Training Data Selection for Efficient Pruned Large Language Model Recovery

### v2

@@ Appendix A Supporting Materials for Introduction Part
Figure 3: Average performance on seven common LLM reasoning evaluation tasks after recovery post-training with different data. The numbers in brackets represent the group index of the data subset in the full dataset. Unpruned indicates the original model and w/o Training indicates the pruned model (using LLM-Pruner [Ma et al., 2023]) without the recovery post-training.
[…]
@@ Appendix D Evaluation on Mathematical Problem Solving Tasks
  • GSM8K [Cobbe et al., 2021]: A dataset containing 8.5K high-quality grade school math word problems that test various mathematical problem solving capabilities, including arithmetic, algebra, and word problem solving.
  • Minerva Math [Lewkowycz et al., 2022]: A comprehensive mathematical evaluation dataset covering diverse topics in mathematics ranging from arithmetic to calculus, with problems requiring multi-step reasoning.
@@ Appendix E Evaluation on Code Generation Tasks
  • HumanEval [Chen et al., 2021]: A widely-used code generation benchmark consisting of 164 hand-crafted Python programming problems that test various programming concepts. Each problem contains a function signature with docstring and test cases, requiring models to complete the implementation. The benchmark evaluates functional correctness using the Pass@$k$ metric, which measures the percentage of problems where a correct solution appears in k samples.
  • MBPP [Austin et al., 2021]: A programming benchmark containing 974 Python programming problems focused on basic programming tasks. Each problem includes a natural language description, test cases, and a reference solution, making it particularly suitable for evaluating language-to-code generation capabilities. MBPP uses the same Pass@$k$ evaluation metric as HumanEval but generally features simpler problems with a broader coverage of basic programming concepts.
@@
Table 7: Recovery performance of different instruction tuning data selection methods under various LLM quantization schemes on LLaMA2-7B model. The ‘bold’ represents the best performance under the same quantization scheme. Here, the Alpaca is taken as the original dataset.
@@
Table 8: Recovery performance of different instruction tuning data selection methods under various LLM quantization schemes on LLaMA2-13B model. The ‘bold’ represents the best performance under the same quantization scheme. Here, the Alpaca is taken as the original dataset.
@@
Appendix F Extended Experiments on Post-quantization Recovery Training
Though the method descriptions and the experiments in the main body are mainly around the LLM pruning scenario, our PASER framework can actually be extended seamlessly to other LLM compression scenario. To further demonstrate its applicability, we conduct the experiments on post-quantization recovery training and compare our PASER with corresponding instruction tuning data selection baselines. In detail, we choose two most representative methods: Round-To-Nearest (RTN) [Frantar and Alistarh, 2022, Yao et al., 2022], GPTQ [Frantar et al., 2023], to perform the LLM quantization. It should be clarified that RTN method, which rounds all weights to the nearest quantized value on exactly the same asymmetric per-row grid, is actually the fundamental technique in most works about LLM quantization [Frantar and Alistarh, 2022, Yao et al., 2022, Park et al., ]. Its runtimes scales well to the models with many billion parameters due to the direct rounding. According to the results provided in Table 7 and 8, we can observe that the PASER can still effectively enhance the recovery performance and outperform the data selection baselines on averaged reasoning performance and zero-shot perplexity for both LLaMA2-7B and LLaMA2-13B models. Meanwhile, recovery data selection baselines can indeed achieve the stronger performance than full data and random baselines, which validates the necessity of conducting recovery data selection even in the LLM quantization scenario. Furthermore, comparing these results with Table 1 and 2, it can be noticed that the improvement space of PASER in Table 7 and 8 has been reduced to some extent. This is because the post-compression performance of such quantization schemes has been competitive enough, which can reflected from the w/o training row.
@@ Appendix F Extended Experiments on Post-quantization Recovery Training
Appendix G Extended Experiments on Recovery Training with Knowledge Distillation
Inspired by  [Muralidharan et al., 2024], we explore the knowledge distillation as the recovery post-training paradigm instead of the standard supervised learning with the groundtruth label. Here, we set the original model $M_{o}$ as the teacher and the pruned model $M_{p}$ as the student. The mean KL divergence [Kullback and Leibler, 1951] between the output probability distribution of $M_{o}$ and that of $M_{p}$ is taken as the objective function. Comparing the corresponding results under different pruning schemes in Table 9 with that in Table 1, we can first observe that knowledge distillation can effectively improve the recovery performance on both reasoning and language modeling tasks in most cases. In particular, the reasoning performance of PASER is improved by 0.348 points on average among such four pruning schemes. Interestingly, the knowledge distillation recovery performance of Full Data under LLM-Pruner is much better than that with standard label-supervised learning. This demonstrates that knowledge distillation is also a promising approach to avoid the misleading information from the irrelevant or conflicting samples existing in the original dataset. Because its learning process directly imitates the unpruned model behavior instead of the provided labels, thus better preserving the thinking and decision-making consistency with the original model. As a summary, distilling the knowledge of unpruned model into the pruned model can be regarded as an effective way to enhance the recovery performance, though bring more memory overhead. Furthermore, stronger layer-wise distillation can also be taken into consideration [Jiao et al., 2020].
@@ Appendix G Extended Experiments on Recovery Training with Knowledge Distillation
Appendix H Experiments on More Recent LLMs
In addition to the previous experiments conducted on LLaMA2/3 and Baichuan2 families, we also supplement some results on more recent and advanced LLMs: LLaMA3.1-8B  [^2] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen2.5-7B [Yang et al., 2024]  [^3] https://huggingface.co/Qwen/Qwen2.5-7B, Qwen3-8B  [^4] Qwen/Qwen3-8B. Besides, the Mixture of Experts(MOE) model Mixtral-8$\times$7B [Jiang et al., 2024]  [^5] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 is also employed to conduct the experiments. The evaluation results have been provided in Table 11. From Table 11, we can observe that PASER consistently outperforms all baseline methods across different recent and advanced LLMs, including the MOE architecture (Mixtral-8x7B). For LLaMA3.1-8B, PASER achieves perplexity scores (7.56 on WikiText2, 17.83 on PTB) that are remarkably close to those of the unpruned model (6.63 and 16.72, respectively), with only a 0.76 point gap in reasoning performance (70.05 vs. 70.81). Notably, PASER demonstrates robust performance across different model architectures. For the MoE-based Mixtral-8x7B, PASER significantly narrows the performance gap from 10.48 points (w/o Training: 61.56 vs. w/o pruning: 72.04) to just 1.17 points (PASER: 70.87 vs. w/o pruning: 72.04), representing a 88.8% recovery of lost capability. This demonstrates that our capability-aware selection approach effectively addresses the specialized recovery needs of sparse models with expert-based architectures. Among the most recent models, Qwen3-8B shows the strongest recovery with PASER, achieving 98.4% of the unpruned model’s reasoning performance (70.58 vs. 71.73). Meanwhile, Qwen2.5-7B demonstrates slightly lower recovery effectiveness, potentially due to differences in its architecture and pretraining approach. These results further validate PASER’s applicability across diverse and state-of-the-art LLM architectures, confirming that our capability degradation-aware instruction selection and negative tuning effects mitigation strategies generalize well beyond the models presented in the main paper.
@@ Appendix H Experiments on More Recent LLMs
Appendix I Detailed Ablation Study Results
@@
Appendix J Hyperparameter Robustness Analysis
@@
J.1 Sensitivity to Embedding Dimension
The dimensionality of the manifold representation after applying the diffusion kernel is controlled by parameter $d$. We evaluate PASER’s performance with $d$ ranging from 4 to 64 on LLaMA2-7B under LLM-Pruner (ratio=25%). As shown in left subfigure of Figure 5, clustering results remains consistent (Rand Index $>$0.85) across this range, and average reasoning performance variations is controlled within $\pm$0.2 points. This indicates that PASER is relatively stable to the specific choice of embedding dimension, though we select $d=16$ based on the eigenvalue decay pattern, where eigenvalues beyond this dimension contribute negligibly to the representation.
@@
J.2 Robustness to Temperature Parameter
For the Jensen-Shannon divergence (JSD) calculation in the capability degradation score, we investigate the effect of temperature $\tau$ in the softmax function (Equation 5). In right subfigure of Figure 5, we test values in the range $\tau\in\{0.5,0.75,1.0,1.25,1.5\}$ and observe that while lower temperatures ($\tau<0.75$) lead to slightly sharper capability distinctions, the overall performance differences are minimal ($<$0.3 points). Our default setting of $\tau=1.0$ provides a good balance between sensitivity to capability differences and stability across pruning schemes.
@@
Appendix K Empirical Exploration for Component Selection
@@
K.1 Exploration on Other Possible Instruction Embedding Models
In the Semantic-Structal Recovery Instruction Clustering module of our PASER framework, we adopt the SentenceBERT [Reimers and Gurevych, 2019] to obtain the initial instruction embedding, considering that SentenceBERT has demonstrated strong transfer capability across various text semantic tasks. To further validate that SentenceBERT is competitive enough here with the empirical evidence, we replace it with two other advanced LLMs: Qwen2.5-7B [Yang et al., 2024] and LLaMA3-8B [Dubey et al., 2024] here and conduct the experiments. As shown in Table 13, using much larger pretrained language models like Qwen2.5-7B and LLaMA3-8B as embedding generators provides marginal performance improvements (less than 0.05 points on average) compared to SentenceBERT. While LLaMA3-8B achieves the highest performance in three configurations, these gains are negligible given the substantially increased computational costs associated with using such large models for initial embeddings. Besides, under the SparseGPT, sparsity=50% setting, SentenceBERT performs obviously better than LLaMA3-8B and Qwen2.5-7B. This finding validates our design choice of SentenceBERT as the default embedding model for PASER, as it strikes an optimal balance between semantic representation quality and computational efficiency. SentenceBERT requires significantly fewer resources while still enabling effective capability-specific clustering that supports our targeted recovery approach. The results demonstrate that while more powerful embedding models may theoretically provide better semantic representations, in practice, the performance difference is minimal for our specific task of recovery data selection.
@@
K.2 Exploration on Other Possible Dimensionality Reduction Methods
@@
K.3 Exploration on Other Possible Clustering Methods
To discuss the impact of different instruction tuning data clustering approaches, we replace our Semantic-structural Recovery Instruction Clustering (S2RIC) module with some other common text clustering method: NMF_TFIDF, LDA_TFIDF, KMeans_TFIDF, Spectral_MTEB, Spectral_BERT [Xu et al., 2024a]. The reasoning performance comparison among different PASER versions with such clustering methods is provided in Table 15. From the table, we can find that integrating other instruction clustering methods with PASER can bring the performance decline to some extent among all four pruning schemes. Especially, the clustering method with traditional statistics-based text representation technique, TFIDF, generally behaves worse than semantic embedding-based text representation techniques like BERT. Therefore, we can conclude that our semantic-structural recovery instruction clustering is at least a competitive approach as the clustering component of PASER. Though, comparing these results with those in Table 1, we can observe the advantages of PASER over other general instruction tuning data selection methods can still be stably maintained. This further demonstrates that the potential of the clustering-based data selection for effective and balanced LLM capability recovery.

### v3

@@ 5 Conclusion and Future Works
Ethics Statement
Reproducibility Statement
We provide a code repository https://github.com/BokwaiHo/PASER containing the implementation of PASER, along with detailed instructions for reproducing our experiments. All datasets used in this work, including Alpaca and LaMini for instruction tuning and standard reasoning/QA benchmarks (e.g., BoolQ, PIQA, ARC, HellaSwag, WinoGrande, OpenbookQA), are publicly available. Hyperparameters, pruning schemes, and evaluation protocols are described in the main text and appendices to ensure clarity. Our theoretical analysis, ablation studies, and efficiency measurements are also documented to facilitate independent verification
Acknowledgments
This work was supported by the Early Career Scheme (No. CityU 21219323) and the General Research Fund (No. CityU 11220324) of the University Grants Committee (UGC), the NSFC Young Scientists Fund (No. 9240127), and the Donation for Research Projects (No. 9229164 and No. 9220187).
@@ Appendix A Supporting Materials for Introduction Part
Figure 3: Average performance on seven common LLM reasoning evaluation tasks after recovery post-training with different data. The numbers in brackets represent the group index of the data subset in the full dataset. Unpruned indicates the original model and w/o Training indicates the pruned model (using LLM-Pruner (Ma et al., 2023)) without the recovery post-training.
Figure 4: Normalized performance degradation degree($\%$) on four various capabilities under four LLM pruning settings.
@@ Appendix D Evaluation on Mathematical Problem Solving Tasks
  • GSM8K (Cobbe et al., 2021): A dataset containing 8.5K high-quality grade school math word problems that test various mathematical problem solving capabilities, including arithmetic, algebra, and word problem solving.
  • Minerva Math (Lewkowycz et al., 2022): A comprehensive mathematical evaluation dataset covering diverse topics in mathematics ranging from arithmetic to calculus, with problems requiring multi-step reasoning.
@@ Appendix E Evaluation on Code Generation Tasks
  • HumanEval (Chen et al., 2021): A widely-used code generation benchmark consisting of 164 hand-crafted Python programming problems that test various programming concepts. Each problem contains a function signature with docstring and test cases, requiring models to complete the implementation. The benchmark evaluates functional correctness using the Pass@$k$ metric, which measures the percentage of problems where a correct solution appears in k samples.
  • MBPP (Austin et al., 2021): A programming benchmark containing 974 Python programming problems focused on basic programming tasks. Each problem includes a natural language description, test cases, and a reference solution, making it particularly suitable for evaluating language-to-code generation capabilities. MBPP uses the same Pass@$k$ evaluation metric as HumanEval but generally features simpler problems with a broader coverage of basic programming concepts.
@@
Table 7: Recovery performance of different instruction tuning data selection methods on other knowledge extensive tasks under different LLM pruning schemes. The ‘bold’ represents the best performance under the same pruning scheme. Here, the Alpaca is taken as the original dataset. The LLaMA2-7B is taken as the target here to ensure consistency.
Benchmark Pruning w/o Training Full Data Random Instruction Mining IFD Nuggets PASER
MMLU LLM-Pruner (25%) $\downarrow$ 25.35 36.27 40.91 41.42 41.86 42.30 46.37
SliceGPT (25%) 28.80 40.82 41.76 42.15 42.51 42.64 46.97
Wanda (2:4) 32.58 44.45 42.36 42.53 42.73 43.55 47.04
SparseGPT (50%) 34.13 45.01 42.47 44.19 43.63 43.18 47.43
CommonsenseQA LLM-Pruner (25%) $\downarrow$ 48.47 57.50 60.21 60.65 61.22 62.59 66.44
SliceGPT (25%) 49.60 61.32 61.09 60.81 62.28 62.23 66.51
Wanda (2:4) 52.65 65.04 61.71 62.13 62.47 62.91 66.79
SparseGPT (50%) 53.97 64.32 62.21 63.04 62.48 63.17 67.63
Appendix F Evaluation on Other Knowledge-extensive Tasks
To supplement more experiments on other knowledge-extensive evaluation tasks, we conduct experiments on MMLU (Hendrycks et al., ) and CommonsenseQA (Talmor et al., 2019), with LLaMA2-7B as the target LLM to ensure comparison consistency. As shown in the Table 7, PASER consistently outperforms existing data selection baselines and even surpasses full-data tuning. This demonstrates that PASER generalizes well beyond the commonsense reasoning benchmarks in the main text and is effective on other knowledge-intensive tasks.
Table 8: Recovery performance of different instruction tuning data selection methods under RTN 4bits and GPTQ 4bits schemes on LLaMA2-7B model. The ‘bold’ represents the best performance under the same quantization scheme. Here, the Alpaca is taken as the original dataset.
@@ Appendix F Evaluation on Other Knowledge-extensive Tasks
Table 9: Recovery performance of different recovery data selection methods under RTN 4bits and GPTQ 4bits quantization schemes on LLaMA2-13B model. The ‘bold’ represents the best performance under the same quantization scheme. Here, the Alpaca is taken as the original dataset.
@@
Table 10: Recovery performance of different recovery data selection methods on LLaMA2-7B under AWQ 4bits scheme. The ‘bold’ represents the best performance among different selection methods. Here, the Alpaca is taken as the original dataset. The “Reason” indicates the averaged performance on 7 commonsense reasoning datasets.
Quantization Benchmark w/o Training Full Data Random Instruction Mining IFD Nuggets PASER
[…]
Appendix G Extended Experiments on Post-quantization Recovery Training
Though the method descriptions and the experiments in the main body are mainly around the LLM pruning scenario, our PASER framework can actually be extended seamlessly to other LLM compression scenario. To further demonstrate its applicability, we conduct the experiments on post-quantization recovery training and compare our PASER with corresponding instruction tuning data selection baselines. In detail, we choose two most representative methods: Round-To-Nearest (RTN) (Frantar & Alistarh, 2022; Yao et al., 2022), GPTQ (Frantar et al., 2023) to perform the LLM quantization. It should be clarified that RTN method, which rounds all weights to the nearest quantized value on exactly the same asymmetric per-row grid, is actually the fundamental technique in most works about LLM quantization (Frantar & Alistarh, 2022; Yao et al., 2022; Park et al., ). Its runtimes scales well to the models with many billion parameters due to the direct rounding. According to the results provided in Table 8 and 9, we can observe that the PASER can still effectively enhance the recovery performance and outperform the data selection baselines on averaged reasoning performance and zero-shot perplexity for both LLaMA2-7B and LLaMA2-13B models. Meanwhile, recovery data selection baselines can indeed achieve the stronger performance than full data and random baselines, which validates the necessity of conducting recovery data selection even in the LLM quantization scenario. Furthermore, comparing these results with Table 1 and 2, it can be noticed that the improvement space of PASER in Table 8 and 9 has been reduced to some extent. This is because the post-compression performance of such quantization schemes has been competitive enough, which can reflected from the w/o training row. In addition, we also conduct experiments on one of the most robust LLM quantization scheme, AWQ (Lin et al., 2024), and the results are provided in the Table 10. Considering the superior performance of AWQ itself on reducing the performance drop brought by the precision loss, our improvement over baselines is still obvious, further demonstrating the generalizability of PASER on quantization schemes.
Table 11: Recovery performance of different instruction tuning data selection methods on LLaMA2-7B under different LLM distillation schemes. The ‘bold’ represents the best performance on the same distillation scheme. Here, the Alpaca is taken as the original dataset. The “Reason” indicates the averaged performance on 7 commonsense reasoning datasets.
Distillation Benchmark w/o Training Full Data Random Instruction Mining IFD Nuggets PASER
[…]
Appendix H Extended Experiments on Post-distillation Recovery Training
In addition to the post-pruning and post-quantization experiments, we also supplement post-distillation experiments to achieve a full LLM post-compression recovery training picture. In detail, we adopt DistiLLM (Ko et al., 2024), MiniPLM (Gu et al., ), SKD (Xu et al., ), respectively, to distill a LLaMA2-7B model into a 1.5B scale small language model. The corresponding post-distillation recovery performance has been provided in the Table 11. From the table, we can observe that our PASER can effectively help recover LLM capabilities compressed with different recent distillation methods. Therefore, we can conclude that our proposed PASER can generalize well across different kinds of LLM compression paradigms, not limited to the pruning discussed in the main text.
@@ Appendix H Extended Experiments on Post-distillation Recovery Training
Appendix I Experiments on Recovery Training with Knowledge Distillation
Inspired by  (Muralidharan et al., 2024), we explore the knowledge distillation as the recovery post-training paradigm instead of the standard supervised learning with the groundtruth label. Here, we set the original model $M_{o}$ as the teacher and the pruned model $M_{p}$ as the student. The mean KL divergence (Kullback & Leibler, 1951) between the output probability distribution of $M_{o}$ and that of $M_{p}$ is taken as the objective function. Comparing the corresponding results under different pruning schemes in Table 12 with that in Table 1, we can first observe that knowledge distillation can effectively improve the recovery performance on both reasoning and language modeling tasks in most cases. In particular, the reasoning performance of PASER is improved by 0.348 points on average among such four pruning schemes. Interestingly, the knowledge distillation recovery performance of Full Data under LLM-Pruner is much better than that with standard label-supervised learning. This demonstrates that knowledge distillation is also a promising approach to avoid the misleading information from the irrelevant or conflicting samples existing in the original dataset. Because its learning process directly imitates the unpruned model behavior instead of the provided labels, thus better preserving the thinking and decision-making consistency with the original model. As a summary, distilling the knowledge of unpruned model into the pruned model can be regarded as an effective way to enhance the recovery performance, though bring more memory overhead. Furthermore, stronger layer-wise distillation can also be taken into consideration (Jiao et al., 2020).
@@ Appendix I Experiments on Recovery Training with Knowledge Distillation
[…]
Appendix J Experiments on More Recent LLMs
[…]
Appendix K Discussion on Comparison with General Data Selection Methods
To further demonstrate our PASER is an unique contribution and solution for pruned-LLM recovery, we discuss the comparison with previous general data selection methods (e.g., LESS (Xia et al., 2024b), DELIFT (Agarwal et al., 2025), SMART (Renduchintala et al., 2024), Datamodels (Ilyas et al., 2022)) from four perspectives: 1) Why general data selection methods do not suffice for pruned-LLM recovery? (Appendix K.1) 2) How PASER differs from previous general data selection methods? (Appendix K.2) 3) What is specific comparison between PASER and gradient-based sampling? (Appendix K.3) 4) Can PASER benefit from gradient-derived dense information? (Appendix K.4)
K.1 Q1: Why General Data Selection Methods do not Suffice for Pruned-LLM Recovery?
While existing data selection methods such as LESS (Xia et al., 2024b), DELIFT (Agarwal et al., 2025), SMART (Renduchintala et al., 2024), and Datamodels (Ilyas et al., 2022) have demonstrated effectiveness for various training objectives (e.g., improving generalization, robustness, or efficiency), they are not specifically tailored to the unique recovery challenges posed by pruning. Notably:
  • Uneven Capability Degradation: Pruning introduces non-uniform degradation across different model capabilities (e.g., commonsense, coding, math), as shown in Figure 4 of Appendix A. General selection methods typically assume uniform importance across examples or rely on training dynamics in full-capacity models, failing to prioritize capability-specific weaknesses that are critical in post-pruning recovery (see Section 1 and Section 3.3). For example, as discussed in the Line 48-50, existing data selection methods generally favor samples with clear structure and natural, human-like expressions. However, these samples may not exactly target specific capabilities (e.g., math, code) severely compromised during pruning. In fact, in the post-pruning recovery scenario, even if the expression of samples regarding severely compromised capabilities is not so clear or structured, we should still prioritize them.
  • Performance Sensitivity to Instruction Content: As shown empirically in Figure 3 of Appendix A, recovery performance varies drastically depending on which subset is selected, an issue rarely addressed by general-purpose selection algorithms. This is a unique challenge in pruned model recovery, as pruning disrupts the model’s internal structure and weakens its representational capacity, making the choice of recovery data more critical and sensitive than in standard fine-tuning scenarios.
  • Incompatibility with Recovery Training Goals: Many existing methods select samples that maximize representativeness or informativeness during pre-training or fine-tuning, which may inadvertently include instructions that are irrelevant or conflicting with each other, thus resulting in harmful impact for recovering pruned capabilities.
Thus, applying these methods directly without modification often leads to inefficient use of recovery budgets and fails to address the nuanced degradation patterns introduced by pruning.
K.2 Q2: How PASER Differs from Previous General Data Selection Methods?
PASER is designed from first principles for post-pruning recovery and introduces three key innovations absent in prior methods:
  • Capability-Aware Clustering and Budget Allocation: Unlike data selection methods that focus on sample diversity or informativeness globally, PASER performs semantic-structural clustering (Section 3.2) to uncover latent instruction groups aligned with distinct capabilities. We then assess degradation in each cluster using Jensen-Shannon Divergence between pruned and original model outputs (Section 3.3), and allocate recovery budget proportionally. This enables targeted and balanced recovery, which existing methods do not consider. In contrast, existing methods typically treat all capabilities uniformly and are agnostic to which skills the model has lost post-pruning. As a result, they may waste recovery budget on less-affected areas while failing to restore critical capabilities.
  • Efficiency-Driven Sample Selection: PASER integrates computational efficiency by computing a sample-wise efficiency score based on degradation severity and expected training cost. This is crucial in recovery settings where data budgets and compute are highly constrained. In contrast, datamodel-based selection methods (e.g., Datamodels (Ilyas et al., 2022)) require training large probes or repeated forward/backward passes, making them computationally expensive and impractical for efficient recovery. LESS is also time-consuming, because it involves gradient computation and LoRA training during the selection process.
  • Negative Tuning Mitigation via Concept Consistency Graph: General-purpose selection approaches rarely guard against harmful instructions. PASER introduces a Concept Consistency Graph (Section 3.4) to detect and filter irrelevant or conflicting samples that might lead to negative tuning effects. This is a particularly pressing issue when the model’s representational capacity has been impaired by pruning, which results in models more sensitive to conflicting information. However, existing methods rarely address this risk, most of which assume all data in high-quality pools is beneficial. However, our results (e.g., Table 1) show that even using the full dataset can degrade performance due to negative tuning effects, especially for pruned models with weakened internal representations
Besides, we would like to highlight that across all pruning schemes and models (Tables 1, 2, 3), PASER consistently outperforms random, full data, and several instruction-tuning selection baselines (e.g., Nuggets, IFD), despite using only 4%–20% of the original data. Additionally, our ablation study (Table 4) confirms the importance of each PASER module in achieving superior recovery.
In summary, PASER is motivated by and explicitly designed for the unique characteristics and challenges of pruned LLM recovery. Existing general data selection methods neither target uneven capability loss nor mitigate the risk of irrelevant or conflicting data. PASER fills this gap by introducing capability-driven, efficiency-aware, and consistency-preserving data selection, thus leading to substantial gains in both recovery quality and training cost.
K.3 Q3: What is Specific Difference between PASER and Gradient-based Sampling?
We would like to clarify the fundamental differences between PASER and gradient-based data selection strategies such as LESS (Xia et al., 2024b), and supplement this with a controlled empirical comparison.
Conceptual Differences Gradient-based methods (e.g., LESS) typically rely on computing influence scores or gradient norms to identify impactful training samples. These scores are often computed with respect to specific layers or objectives (e.g., classification loss or instruction adherence) and require gradient backpropagation, which incurs high computational cost—especially when evaluating all samples across large instruction tuning sets. In contrast, PASER is specifically designed for capability recovery after structured/semi-structured/unstructured pruning, where the pruned model exhibits uneven degradation across semantic capabilities. Rather than using gradient signals, PASER leverages: 1) Semantic clustering to disentangle instruction groups aligned with distinct capabilities; 2) Capability Degradation Scores (CDS) based on Jensen-Shannon divergence (JSD) between the pruned and original models’ output distributions; 3) Efficiency-aware scoring to optimize recovery benefit per unit of compute cost; 4) Negative tuning effect mitigation using concept consistency graphs. This design enables PASER to approximate dense information signals (through token-level divergence) while being more effective, efficient, and robust than backpropagation-dependent strategies.
Empirical Comparison To further clarify the performance and efficiency differences, we compare PASER with a LESS-style gradient-based sampling baseline, which selects the top 20% samples from Alpaca with the highest gradient norm of loss w.r.t. embedding parameters.
Table 15: Recovery performance of different data selection methods on LLaMA2-7B under different LLM distillation schemes. Here, the Alpaca is taken as the original dataset. The “Reason” indicates the averaged performance on 7 commonsense reasoning datasets.
Pruning
Recovery
Data Selection
Reasoning $\uparrow$ WikiText2 $\downarrow$ PTB $\downarrow$
Data Selection
Time (mins) $\downarrow$
Recovery Training
Time (mins) $\downarrow$
[…]
As shown in Table 15, PASER not only achieves better recovery performance across both language modeling and reasoning tasks, but also incurs significantly less data selection time (up to 30× less) due to its gradient-free design. The recovery training time is also reduced ($\downarrow$  14.2%), as PASER’s efficiency-driven sampling strategy favors shorter and semantically cohesive instructions. Furthermore, PASER consistently exhibits lower performance variance across different pruning settings (e.g., reasoning accuracy std of 0.08–0.14 vs. 0.35–0.51 for LESS-style), indicating greater robustness under heterogeneous post-pruning degradation. In summary, PASER offers a more effective, efficient, robust, and pruning-aware alternative to gradient-based sampling strategies, especially in scenarios where gradient signals become noisy or unreliable due to structural capacity loss.
K.4 Q4: Can PASER Benefit from Gradient-derived Dense Information?
While PASER does not explicitly rely on gradient information, we acknowledge that gradient-derived signals capture dense local sensitivity and can, in principle, provide useful guidance for data selection. In our current design, we approximate similar effects using token-level output distributional divergence (JSD) to reflect fine-grained capability deterioration without requiring backward computation. This makes PASER highly efficient and model-agnostic. That said, we agree with the reviewer that incorporating lightweight gradient signals could further enhance selection granularity. In future work, we plan to explore hybrid approaches by integrating gradient-based information into PASER in the following ways:
  • Gradient-Guided Cluster Weighting: Use average gradient norms within each capability cluster to refine degradation scores alongside our JSD-based measure.
  • Gradient-Augmented Efficiency Scoring: Combine gradient-based sample importance with our current JSD-based efficiency metric, possibly as a weighted fusion.
  • LoRA-style Gradient Probing: Introduce low-rank adapters to cheaply approximate gradient information in a parameter-efficient manner during selection, especially useful for large LLMs.
We believe this line of integration can combine the dense expressiveness of gradients with PASER’s efficient, capability-aware framework, further improving recovery quality without incurring excessive compute overhead.
@@
Appendix L Detailed Ablation Study Results
@@
Appendix M Hyperparameter Robustness Analysis
@@
M.1 Sensitivity to Embedding Dimension
The dimensionality of the manifold representation after applying the diffusion kernel is controlled by parameter $d$. We evaluate PASER’s performance with $d$ ranging from 4 to 64 on LLaMA2-7B under LLM-Pruner (ratio=25%). As shown in left subfigure of Figure 5, clustering results remains consistent (Rand Index >>0.85) across this range, and average reasoning performance variations is controlled within $\pm$0.2 points. This indicates that PASER is relatively stable to the specific choice of embedding dimension, though we select $d=16$ based on the eigenvalue decay pattern, where eigenvalues beyond this dimension contribute negligibly to the representation.
@@
M.2 Robustness to Temperature Parameter
For the Jensen-Shannon divergence (JSD) calculation in the capability degradation score, we investigate the effect of temperature $\tau$ in the softmax function (Equation 5). In right subfigure of Figure 5, we test values in the range $\tau\in\{0.5,0.75,1.0,1.25,1.5\}$ and observe that while lower temperatures ($\tau<0.75$) lead to slightly sharper capability distinctions, the overall performance differences are minimal (<<0.3 points). Our default setting of $\tau=1.0$ provides a good balance between sensitivity to capability differences and stability across pruning schemes.
@@
Appendix N Empirical Exploration for Component Selection
@@
N.1 Exploration on Other Possible Instruction Embedding Models
In the Semantic-Structal Recovery Instruction Clustering module of our PASER framework, we adopt the SentenceBERT (Reimers & Gurevych, 2019) to obtain the initial instruction embedding, considering that SentenceBERT has demonstrated strong transfer capability across various text semantic tasks. To further validate that SentenceBERT is competitive enough here with the empirical evidence, we replace it with two other advanced LLMs: Qwen2.5-7B (Yang et al., 2024) and LLaMA3-8B (Dubey et al., 2024) here and conduct the experiments. As shown in Table 17, using much larger pretrained language models like Qwen2.5-7B and LLaMA3-8B as embedding generators provides marginal performance improvements (less than 0.05 points on average) compared to SentenceBERT. While LLaMA3-8B achieves the highest performance in three configurations, these gains are negligible given the substantially increased computational costs associated with using such large models for initial embeddings. Besides, under the SparseGPT, sparsity=50% setting, SentenceBERT performs obviously better than LLaMA3-8B and Qwen2.5-7B. This finding validates our design choice of SentenceBERT as the default embedding model for PASER, as it strikes an optimal balance between semantic representation quality and computational efficiency. SentenceBERT requires significantly fewer resources while still enabling effective capability-specific clustering that supports our targeted recovery approach. The results demonstrate that while more powerful embedding models may theoretically provide better semantic representations, in practice, the performance difference is minimal for our specific task of recovery data selection.
@@
N.2 Exploration on Other Possible Dimensionality Reduction Methods
@@
N.3 Exploration on Other Possible Clustering Methods
To discuss the impact of different instruction tuning data clustering approaches, we replace our Semantic-structural Recovery Instruction Clustering (S2RIC) module with some other common text clustering method: NMF_TFIDF, LDA_TFIDF, KMeans_TFIDF, Spectral_MTEB, Spectral_BERT (Xu et al., 2024a). The reasoning performance comparison among different PASER versions with such clustering methods is provided in Table 19. From the table, we can find that integrating other instruction clustering methods with PASER can bring the performance decline to some extent among all four pruning schemes. Especially, the clustering method with traditional statistics-based text representation technique, TFIDF, generally behaves worse than semantic embedding-based text representation techniques like BERT. Therefore, we can conclude that our semantic-structural recovery instruction clustering is at least a competitive approach as the clustering component of PASER. Though, comparing these results with those in Table 1, we can observe the advantages of PASER over other general instruction tuning data selection methods can still be stably maintained. This further demonstrates that the potential of the clustering-based data selection for effective and balanced LLM capability recovery.